OpenStudy (anonymous):

what is the derivative of y=cos^5[pi(t)-19]

myininaya (myininaya):

\[y=[\cos(\pi t -19)]^5\]

myininaya (myininaya):

you have a lot of chain rule that needs to happen

myininaya (myininaya):

\[y'=5 [\cos(\pi t -19)]^{5-1}[\cos(\pi t -19)]' \text{ by chain rule and power rule }\]

myininaya (myininaya):

see if you can continue from here